40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 11 15:04:52 2009, Alberto, Update, ABSL, Working on the AP table 
    Reply  Wed Nov 11 16:18:57 2009, Alberto, Update, ABSL, Working on the AP table 
       Reply  Wed Nov 11 17:10:51 2009, Jenne, Update, ABSL, Working on the AP table 
       Reply  Wed Nov 11 17:33:54 2009, Koji, Update, ABSL, Working on the AP table 
          Reply  Thu Nov 12 11:34:38 2009, Alberto, Update, ABSL, Working on the AP table 
             Reply  Thu Nov 12 16:53:59 2009, Alberto, Update, ABSL, Working on the AP table 
Message ID: 2252     Entry time: Thu Nov 12 11:34:38 2009     In reply to: 2241     Reply to this: 2257
Author: Alberto 
Type: Update 
Category: ABSL 
Subject: Working on the AP table 

Quote:

Yes it did.

For long time, the crystal temperature C1:PSL-126MOPA_LTMP was 43~46deg. Now it is 34deg. Try ~10deg lower temperature.

Quote:

I wonder if the all the tinkering on the PSL laser done recently to revive the power have changed the PSL NPRO temperature and so its frequency. That could also explain why the beat doesn't show up at the same temperature of the NPRO as I used to operate it. Although I scanned the NPRO temperature +/- 2 deg and didn't see the beat.

 

 Beat found at 30MHz with auxiliary NPRO temperature of 37.19 degrees, vs. ~48 deg as it used to be.

The beat is small (-70dBm). PLL alignment has to be improved.

ELOG V3.1.3-